OpenStudy (anonymous):

(-3)(-4)(-2/3)(-1/2)

OpenStudy (anonymous):

just multiply across 4

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

-3 times -4 is ...?

OpenStudy (anonymous):

i know but the fraction part

OpenStudy (anonymous):

ok so you multiply 12 by (-2/3), just multiply the numerators and you get -24/3, do u see how i got that?

OpenStudy (anonymous):

-24/3=-8 -8*-1/2=8/2=4

OpenStudy (anonymous):

yes but how do u combine the fractions and the whole number

OpenStudy (anonymous):

just multiply them

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

do the denominators have to match